OpenStudy (anonymous):

Two planets, planet A and planet B, have the same surface gravity. However, planet B has twice the radius of planet A. How does the mass of planet B compare to the mass of planet A?

OpenStudy (mrnood):

the surface gravity acceleration is given by g = GM/r^2 Where M is the mass of the planet ant r is the radius of the planet If they are the same then GM1/r1^2 = GM2/r2^2 but you know r1=2r2 so you can work out M1/M2